# This work has been submitted to the IEEE Access for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible. xx.xxxx/ACCESS.2024.xxxxxxx

### v2

@@ III-B2 Predictive Coding
This study showed that combination of learning rules is potential for enabling CL. However, this complex combination of multiple learning rules make it challenging for deployments under OCL scenario.
III-B3 Active Dendrites
Studies in [312] proposed an SNN model leveraging active dendrites [313] to facilitate task-specific sub-network formation (Fig. 18). The spike time $t_{j}$ of a neuron is modulated by a function of the selected dendritic segment $u_{jn}$ for the current task. The dendritic activation function $f(u)$ modulates the spike time dynamically, allowing the model to adapt its behavior based on the task context. It also leverages TTFS encoding to introduce a gating mechanism, that enables selection of sub-networks for various tasks. However, this method requires labeled data or supervisory signals, hence making it challenging for deployments under the OCL scenario.
@@ III-B3 Active Dendrites
III-B4 Bayesian Continual Learning
This method represents weights with parameters that quantify the epistemic uncertainty based on prior knowledge and observed data, and employs Bayesian methods for handling uncertainty over time by determining which knowledge to retain and which to forget [297] (Fig. 19). For real-valued synapses, it uses a Gaussian variational distribution to adjust the values; while for binary synapses, it uses a Bernoulli variational distribution with Gumbel-softmax [314]. Although, this method provides better-calibrated decisions and better detection compared to conventional frequentist approaches [315], the uncertainty quantification incurs high computational complexity, thereby making it challenging for deployments in OCL scenario.
@@ III-B4 Bayesian Continual Learning
III-B5 Architecture-based Approach
Dynamic Structure Development of Spiking Neural Networks (DSD-SNN) enhances the SNN structure by growing new neurons for new tasks and pruning redundant neurons [316] (Fig. 20). It employs a deep SNN architecture comprising of multiple convolutional and fully-connected (Fc) layers, which is equipped with random growth, adaptive pruning, and freezing of neuron mechanisms. This study advanced multi-task learning while enhancing memory capacity and efficiency. However, the development of dynamic structure adds complexity to SNN implementation, which is challenging for OCL scenario.
@@ III-B5 Architecture-based Approach
Self-Organized Regulation SNN (SOR-SNN) employs a pathway search module to adaptively activate task-specific sparse neural pathways based on fundamental weights $W_{t}$ [317] (Fig. 21). Each synapse has two states (i.e., active and inactive), and is determined by learnable synaptic selection parameters. The model decides whether to activate or inhibit each weight by comparing the learnable parameter $A_{s}$ with the threshold $\tilde{A}_{s}$, where activation is preferred if $A_{s}>\tilde{A}_{s}$. This study showed that SOR-SNN enables adaptive reorganization, BWT, and self-repair capabilities. Despite its benefits, this method requires complex dynamic connectivity capabilities, hence making it challenging for deployments under OCL scenario.
@@
III-B6 Rehearsal/Replay-based Approach
A memory replay approach using Experience Replay (ER) method has been developed, where a convolutional SNN is trained to learn in Class-IL and TA scenarios [318]. For resource-constrained devices, the memory-efficient Latent Replay (LR)-based method has been developed, which stores compressed data representations [292]. The Latent Replay training involves a pre-training phase, where SNN is divided into two parts i.e., frozen and learning layers (Fig. 22). The network stores latent replays for memory, and only trains the learning layers on new data. Here, the main challenges are related to the efficient store-load mechanisms for replay data under OCL scenario.
Figure 22: SNN with Recurrent Neurons and Latent Replays; adapted from studies in [292]. (a) Recurrent neuron model. (b) Structure of a generic fully-connected model for CL. (c) An example of the data fed to the Latent Replay stage.
III-B7 Regularization-based Approach
@@
This approach employs regularization terms to balance the old and new tasks, such as Noise Regularization [319], Freezing Large Weights, and Stochastic Langevin Dynamics [289]. For instance, Langevin Dynamics exploit the fact that each weight $w_{i}$ can vary without impacting the accuracy in the domain $D_{i}$ if constrained to a specific space when learning new tasks; see Fig. 23. This study highlighted that Langevin dynamics can exhibit very similar performance as the replay-based methods while being less memory-intensive. However, its compute requirements are relatively high, thus requiring further studies for OCL scenario.
III-B8 Hebbian Learning
Hebbian Learning-based Orthogonal Projection (HLOP) method leverages lateral connections and Hebbian learning to achieve CL [320]. It employs orthogonal gradient projection to modify pre-synaptic activity traces (Fig. 24) to ensure that weight updates for new tasks do not interfere with the weights associated with the old tasks. Hebbian learning is useful to extract the principal subspace of neural activities, and updates synaptic weights based on the correlation between pre- and post-synaptic neuron activities, while the activity traces are updated using lateral signals.
@@ III-B8 Hebbian Learning
III-C Efficiency Enhancement Methods for NCL
@@
To reduce the memory footprint and energy consumption of SNNs with CL capabilities (NCL), several optimization methods have been proposed and discussed below.
III-C1 Reduction of SNN Operations
@@ III-C2 Weight Quantization
TABLE IV: Summary of the state-of-the-art works for energy-efficient CL employing bio-plausible SNNs (i.e., NCL).
Work, CL Learning Neuron Neural Learning Learning Optimization Dataset Tools
Year Scenario Setting Model Coding Rule Rate Technique
(2017) [327]
SpikeDyn [41], Class-IL Unsupervised LIF Rate STDP Adaptive Weight Decay MNIST Python
lpSpikeCon [42], Class-IL Unsupervised LIF Rate STDP Adaptive Weight MNIST Python
(2022) Quantization [299]
CFN [96], Class-IL Unsupervised LIF Rate STDP Adaptive - MNIST
SpNCN [296], Task-IL Supervised LIF Rate ST-LRA, - - MNIST [328]
(2023) ST-LRA+ FMNIST [329]
STDP Not-MNIST [330]
[312],(2024) Task-IL Supervised Enhanced TTFS Fixed Weight and Delay Split-MNIST FPGA,
model of Quantization Python

### v3

@@ III-D2 Predictive Coding
A case study was conducted using adapted versions of standard CL benchmarks, i.e., Split MNIST, Split NotMNIST, and Split Fashion MNIST (FMNIST), in a spike-train continuous-time setting. Each dataset was partitioned into five sequential tasks, where each task involved classifying two object categories (e.g., digit pairs in Split MNIST, letters A–J in NotMNIST, or clothing types in FMNIST). The tasks were presented in a fixed sequence, simulating a dynamic CL scenario where the learner must adapt to new tasks while retaining prior knowledge, despite changes in the label distribution that create cross-task interference. Experimental results showed normalized accuracy of 0.9653 on MNIST, 0.9120 on Not-MNIST and 0.9995 on FMNIST (see summary in Table IX and comparative empirical analysis in Table XI and Table XII). This study showed that combination of learning rules is potential for enabling CL. However, this complex combination of multiple learning rules make it challenging for deployments under OCL scenario.
III-D3 Active Dendrites
Studies in [356] proposed an SNN model leveraging active dendrites [357] to facilitate task-specific sub-network formation (Fig. 22). The spike time $t_{j}$ of a neuron is modulated by a function of the selected dendritic segment $u_{jn}$ for the current task. The dendritic activation function $f(u)$ modulates the spike time dynamically, allowing the model to adapt its behavior based on the task context. It also leverages TTFS encoding to introduce a gating mechanism, that enables selection of sub-networks for various tasks. Experimental results on the Split MNIST dataset for sequential CL tasks demonstrated an end-of-training accuracy of 88.3%. Moreover, the FPGA implementation matched the quantized software model with an average inference time of 37.3ms and an accuracy of 80.0%, highlighting the potential for applications in resource-constrained environments (see summary in Table IX and comparative empirical analysis in Table XI and Table XII). However, this method requires labeled data or supervisory signals, hence making it challenging for deployments under the OCL scenario.
@@ III-D3 Active Dendrites
III-D4 Bayesian Continual Learning
This method represents weights with parameters that quantify the epistemic uncertainty based on prior knowledge and observed data, and employs Bayesian methods for handling uncertainty over time by determining which knowledge to retain and which to forget [304] (Fig. 23). For real-valued synapses, it uses a Gaussian variational distribution to adjust the values; while for binary synapses, it uses a Bernoulli variational distribution with Gumbel-softmax [358]. The work was implemented using Intel’s Lava platform [314], enabling Bayesian CL in SNNs, see Table VII. Experimental results achieved avg. test accuracy of $85.44\pm 0.16\%$ with 5× memory reduction on Split-MNIST and 74% for MNIST-DVS (see summary in Table IX and comparative empirical analysis in Table XI and Table XII). Although, this method provides better-calibrated decisions and better detection compared to conventional frequentist approaches [359], the uncertainty quantification incurs high computational complexity, thereby making it challenging for deployments in OCL scenario.
@@ III-D4 Bayesian Continual Learning
III-D5 Architecture-based Approach
Dynamic Structure Development of Spiking Neural Networks (DSD-SNN) enhances the SNN structure by growing new neurons for new tasks and pruning redundant neurons [360] (Fig. 24). It employs a deep SNN architecture comprising of multiple convolutional (CONV) and fully-connected (FC) layers, which is equipped with random growth, adaptive pruning, and freezing of neuron mechanisms. This SNN was implemented using the Brain-inspired Cognitive intelligence engine (BrainCog) [315], see Table VII for details. Experimental results on the MNIST dataset in Task-IL scenario, demonstrated an accuracy of $97.30\pm 0.09\%$ with a network parameter compression rate of 34.38%, outperforming the EWC, GEM and DEN DNN-based CL methods. Furthermore, an accuracy of $96.94\pm 0.05\%$ was reported for the N-MNIST and $77.92\%\pm 0.29$ in Task-IL and 60.47% (10 steps) in Class-IL scenario for the CIFAR100 dataset (see summary in Table IX and comparative empirical analysis in Table XI and Table XII). This study advanced multi-task learning while enhancing memory capacity and efficiency. However, the development of dynamic structure adds complexity to SNN implementation, which is challenging for OCL scenario.
@@ III-D5 Architecture-based Approach
Self-Organized Regulation SNN (SOR-SNN) employs a pathway search module to adaptively activate task-specific sparse neural pathways based on fundamental weights $W_{t}$ [361] (Fig. 25). Each synapse has two states (i.e., active and inactive), and is determined by learnable synaptic selection parameters. The model decides whether to activate or inhibit each weight by comparing the learnable parameter $A_{s}$ with the threshold $\tilde{A}_{s}$, where activation is preferred if $A_{s}>\tilde{A}_{s}$. The experimental results on the CIFAR100 dataset in Class-IL scenario demonstrated an average accuracy of $86.65\%\pm 0.20$ for 20 steps, where each task contains 5 classes and an accuracy improvement of 2.20% compared to the DSD-SNN. Moreover, on the Mini-ImageNet dataset, it reported $>55\%$ accuracy, higher than the existing studies (see summary in Table IX and comparative empirical analysis in Table XI and Table XII). This study showed that SOR-SNN enables adaptive reorganization, BWT, and self-repair capabilities. Despite its benefits, this method requires complex dynamic connectivity capabilities, hence making it challenging for deployments under OCL scenario.
@@
III-D6 Rehearsal/Replay-based Approach
A memory replay approach using ER method has been developed, where a CSNN is trained to learn in Class-IL and TA scenarios [362]. For resource-constrained devices, the memory-efficient Latent Replay (LR)-based method has been developed, which stores compressed data representations [299]. The Latent Replay training involves a pre-training phase, where SNN is divided into two parts i.e., frozen and learning layers. The network stores latent replays for memory, and only trains the learning layers on new data. Here, the main challenges are related to the efficient store-load mechanisms for replay data under OCL scenario. The implementations are based on the Python framework. Experimental results on SHD dataset showcased a Top-1 accuracy of 92% in the CLass-IL scenario and a memory-accuracy trade-off with only 4% accuracy drop due to compression. Unlike prior method [299] that rely on long timesteps and compression-decompression for accuracy, Replay4NCL [363] advances NCL for embedded AI systems by significantly reducing latency and energy consumption. It compressed latent data and replayed it with fewer timesteps during NCL training. Parameter adjustments (e.g., neuron thresholds, learning rate) compensated for reduced spikes. The SNN (Fig. 26) was implemented in Python in a Class-IL scenario with 19 tasks for pre-training and a 20th task for CL. Evaluations on the SHD dataset demonstrated 90.43% accuracy compared to 86.22% of the baseline [299] at layer 3, with 4.88$\times$ speed-up, 20% latent memory saving, and 36.43% energy savings (see summary in Table IX and comparative empirical analysis in Table XI and Table XII).
Figure 26: (a) SNN architecture with recurrent neurons and latent replay (LR) and (b) configuration of current activation data and LR data [363].
III-D7 Regularization-based Approach
This approach employs regularization terms to balance the old and new tasks, such as Noise Regularization [364], Freezing Large Weights, and Stochastic Langevin Dynamics [294]. For instance, Langevin Dynamics exploit the fact that each weight $w_{i}$ can vary without impacting the accuracy in the domain $D_{i}$ if constrained to a specific space when learning new tasks; see Fig. 27. The SNN was implemented in SpykeTorch [312], an open-source PyTorch-based framework supporting non-leaky IF neurons and local learning rules like STDP and R-STDP, see Table VII. The experiments were performed in a Task-IL scenario using the MNIST dataset (i.e., Task1) for initial training in a layer-by-layer training approach, where layer-S1 and layer-S2 were trained using STDP, while the final layer-S3 was trained with R-STDP. The Extended MNIST (EMNIST) dataset containing both letters and digits was then used for subsequent training (i.e., Task 2), assuming MNIST data is not available. The classification accuracy for digits reached $92.0\pm 0.1\%$ and for letters it reached $79.7\pm 0.5\%$ (see summary in Table IX and comparative empirical analysis in Table XI and Table XII). This study highlighted that Langevin dynamics can exhibit very similar performance as the replay-based methods while being less memory-intensive. However, its compute requirements are relatively high, thus requiring further studies for OCL scenario.
@@ III-D7 Regularization-based Approach
III-D8 Hebbian Learning
Hebbian Learning-based Orthogonal Projection (HLOP) method leverages lateral connections and Hebbian learning to achieve CL [365]. It employs orthogonal gradient projection to modify pre-synaptic activity traces (Fig. 28) to ensure that weight updates for new tasks do not interfere with the weights associated with the old tasks. Hebbian learning is useful to extract the principal subspace of neural activities, and updates synaptic weights based on the correlation between pre- and post-synaptic neuron activities, while the activity traces are updated using lateral signals. The implementations are based on the PyTorch framework [319]. Experimental results on various datasets and scenarios, including Permuted MNIST (PMNIST), CIFAR-100, and a combination of datasets like CIFAR-10, MNIST, SVHN, Fashion MNIST, and notM NIST, demonstrated higher accuracy such as 95.15% for PMNIST and lower forgetting (i.e., BWT) across different training methods like Dynamic Spike Representation (DSR), BPTT with SG, and Online Training Through Time (OTTT). HLOP also showcased superior performance compared to DNN-based methods such as EWC, HAT, and GPM highlighting its potential for robust NCL (see summary in Table IX and comparative empirical analysis in Table XI and Table XII).
@@ III-D8 Hebbian Learning
TABLE IX: Summary of the state-of-the-art works for energy-efficient CL employing bio-plausible SNNs (i.e., NCL).
Work, CL Learning Neuron Neural Learning Learning Optimization Dataset Software
Year Scenario Setting Model Coding Rule Rate Technique
(2017) [313]
SpikeDyn [42], Class-IL Unsupervised LIF Rate STDP Adaptive Weight Decay MNIST Bindsnet
(2021)  [306]
lpSpikeCon [43], Class-IL Unsupervised LIF Rate STDP Adaptive Weight MNIST Bindsnet
(2022) Quantization  [306]
CFN [101], Class-IL Unsupervised LIF Rate STDP Adaptive - MNIST -
SpNCN [303], Task-IL Supervised LIF Rate ST-LRA, - - MNIST [366] -
(2023) ST-LRA+ FMNIST [367]
STDP Not-MNIST [368]
[356],(2024) Task-IL Supervised Enhanced TTFS Back- Fixed Weight and Delay Split-MNIST FPGA,
model of Propagation Quantization Python
[369]
Learning MNIST-DVS Lava
[370] [314]
DSD-SNN Task-IL, Supervised LIF - Back- - - MNIST BrainCog
[360],(2023) Class-IL Propagation N-MNIST [340], [315]
Split-CIFAR100
SOR-SNN Class-IL Supervised LIF - Back- Adaptive - CIFAR100 [371], -
[361],(2023) Propagation ImageNet [372]
[362],(2023) Class-IL, Supervised LIF - BPTT Fixed - MNIST PyTorch
TA Fixed [319]
[299],(2023) Sample-IL, LIF - BPTT - Time SHD Python
Class-IL Supervised Compression [343]
Replay4NCL Class-IL Supervised LIF - BPTT Adaptive Timestep SHD Python
 [363],(2025) Reduction [343]
Order R-STDP $a^{-}>-0.1125$ EMNIST [373] [312]
[365],(2024) DIL CIFAR100
III-E Hybrid Learning Paradigms
The scalability of most NCL methods to complicated problems with many tasks or complex inputs is challenging. In this section, we explore hybrid approaches that combine supervised and unsupervised learning paradigms to address CF and improve CL performance. We categorize these approaches into three main classes: 1) self-supervised pre-training hybrids, 2) STDP + supervised learning hybrids, and 3) generative-discriminative hybrid models; see Fig. 29.
Figure 29: A taxonomy of hybrid approaches combining supervised and unsupervised learning paradigms. We have highlighted the main categories (blue blocks), with their works shown (red blocks).
III-E1 Self-supervised Pre-training Hybrid
Cortico-Hippocampal Hybrid Networks (CH-HNN)  [374] integrated ANNs for generalization and SNNs for specific memory encoding, inspired by dual-memory corticohippocampal circuits. It employed ANN-guided episode inference and metaplasticity with SG training to dynamically regulate synaptic plasticity. CH-HNN was validated in two robotic case studies: a Unitree GO1 robot performed real-time MNIST digit recognition using ANN-SNN inference, and a Unitree Z1 arm achieved 82±$\pm 7.25\%$ accuracy in sCIFAR-100 object grasping using YOLO. Across benchmarks like split MNIST, pMNIST, sCIFAR-100, sTiny-ImageNet (70.72% accuracy), and DVS Gesture, CH-HNN outperformed EWC, SI, XdG, iCaRL, and FOSTER in stability-plasticity balance, disparity, and efficiency. For neuromorphic deployment, int8 quantization incurred minimal loss, with SNNs yielding 60.82% power savings over ANNs. Although CH-HNN supports various spiking neuron models, their selection involves trade-offs between computational cost, biological fidelity, and deployment efficiency.
III-E2 STDP + Supervised Learning
The study in [375] integrated ANN-to-SNN conversion with STDB to converge to optimal accuracy in fewer epochs compared to training from scratch. Evaluations on image classification tasks using CIFAR-10, CIFAR-100, and ImageNet for VGG and ResNet architectures showed that SNNs trained with the hybrid method required significantly fewer time steps (10×25×10$\times$-25$\times$10 × - 25 × fewer) to achieve competitive accuracy compared to purely converted SNNs. For instance, the model achieved a top-1 accuracy of 65.19% on the ImageNet dataset using only 250 time steps. Although, the study showed promise, the inherent challenges of converting SNNs, such as their non-differentiable nature, still pose difficulties that may affect broader applicability.
III-E3 Generative-Discriminative Hybrid Models
SpikeGAN [376] is a hybrid generative model that combined an SNN generator with an ANN discriminator to learn and generate temporal spiking data distributions. The SNN generator captured spatio-temporal patterns, while the ANN discriminator enabled adversarial training akin to standard GANs. Bayesian learning was applied to the generator’s weights, and a continual meta-learning framework supported adaptation to multiple real-world distributions. Evaluated on handwritten digit generation, SpikeGAN outperformed ANN classifier accuracy by 20% and enabled SNN classifiers trained on synthetic data to match the performance of those trained on real rate-encoded inputs. The model leveraged the complementary strengths of ANNs and SNNs; however, it may suffer from sample space coverage issues, as indicated by Train-on-Synthetic-Test-on-Real (TSTR) errors, pointing to potential limitations in robustness when generalizing to unseen real data.
Dynamic Lifelong learning with Spiking Generative Networks (DL-SGN) [377] is a hybrid SNN-ANN lifelong learning framework designed to enable energy-efficient image generation and classification on edge devices while addressing CF. It integrated ANN-based dynamic expert modules that grow through Dynamic Knowledge Adversarial Fusion (DKAF), an SNN-based student module combining a VAE and classifier, and an ANN-based assistant discriminator trained adversarially to support generalization. Implemented in PyTorch, DL-SGN achieved substantial improvements in FID scores e.g., MNIST (56.26 vs. 95.26) and CIFAR10 (92.46 vs. 192.2) and surpassed ANN-based replay methods such as LTS and LGM in classification accuracy across complex benchmarks like SVHN-CIFAR10-ImageNet10 task (average: 52.43%). In semi-supervised lifelong learning, it reduces MNIST classification error to 2.47%, significantly outperforming DGR (7.27%). However, the framework incurs notable computational overhead due to expert module expansion and adversarial training. Future directions include optimizing expansion strategies, developing SNN-native generative models, and unifying supervised and unsupervised objectives for improved adaptability in dynamic environments.
Our analysis reveals that hybrid architectures consistently outperform single-paradigm approaches by leveraging complementary strengths of different learning mechanisms. Key observations include the effectiveness of biologically-inspired designs CH-HNN, the scalability advantages of generative replay methods, and the computational efficiency gains from hybrid training protocols. These approaches demonstrate significant improvements in both Task-IL and Class-IL scenarios across standard benchmarks including MNIST, CIFAR-100, and ImageNet variants.
III-F Efficiency Enhancement Methods for NCL
@@
To further reduce the memory footprint and energy consumption of NCL methods, several optimization methods proposed in literature are discussed below, addressing key question Q6.
III-F1 Reduction of SNN Operations
@@ III-F2 Weight Quantization
III-F3 Knowledge Distillation
To enhance the efficiency of SNNs in resource-constrained environments, knowledge distillation has been widely explored as an optimization technique [384, 385, 386, 387]. In particular, teacher-student knowledge distillation enables smaller models to retain the performance of larger models while reducing memory and computational overhead. Shaw et al.  [388] demonstrated the successful application of teacher-student knowledge distillation for radar perception on embedded accelerators, optimizing neural networks for efficient real-time processing. Similar techniques could be leveraged in NCL to improve computational efficiency and reduce model complexity for neuromorphic edge applications.
III-G Compatibility with Neuromorphic Hardware
The requirements for implementing a CL method on neuromorphic hardware include: (1) computational units/modules that can execute the CL algorithm, and (2) sufficient size of computational and memory resources. Specifically, the neuromorphic hardware should have compute units that can execute the given CL algorithm on-chip. For instance, if the CL algorithm employs a bio-plausible STDP learning rule, then the neuromorphic hardware should have a learning unit that can execute it. Meanwhile, the size of computational and memory modules from the neuromorphic hardware defines how much workload can be executed at one time, thereby computational and memory constraints will affect the performance in terms of processing latency, throughput, and energy efficiency.
The compatibility of reviewed methods on the existing neuromorphic hardware can be analyzed based on the capability of the hardware to perform on-chip learning. The summary of existing neuromorphic hardware and their on-chip learning capabilities is provided in Table X. This table shows that, most of neuromorphic hardware platforms do not support on-chip learning, or only support specific on-chip learning mechanisms (e.g., bio-plausible STDP). Therefore, most of reviewed methods may not be executed on the existing neuromorphic hardware, rather they can be executed on the host processing unit (CPU/GPU), which is capable of executing more types of operations, addressing the key question Q7.
TABLE X: Summary of the existing neuromorphic hardware platforms with their on-chip learning capabilities.
Neuromorphic On-chip On-chip Learning
Processor Learning Mechanism
NeuroGrid [325] No -
ROLLS [327] Yes SDSP
TrueNorth [327] No -
SpiNNaker [389] Yes STDP
BrainscaleS-2 [390] Yes STDP, R-STDP, SG
Loihi [328] Yes STDP, Surrogate Gradient Learning
Tianjic [391] No -
MorphIC [392] Yes SDSP
DYNAP [329] No -
Akida [330] Yes Last-layer Learning
PAIBoard [393] No -
PAICORE [394] Yes STDP
TABLE XI: Comparative quantitative analysis of the surveyed NCL methods evaluated on P-MNIST, Split CIFAR-100, Split MNIST, and ImageNet datasets with relevant state-of-the-art DNN-based CL methods addressing energy-efficiency problem with various settings and scenarios.
P-MNIST
Methods Network/ Accuracy CL CL Memory Latency Energy
Model Setting Scenario
GPM [215] FC $93.91\%\pm 0.16$ Supervised Task-IL,
Class-IL
n/a 245s (train) n/a
 [223] FC $84.3\%\pm 0.3$ Supervised Task-IL n/a n/a n/a
A-GEM [222] FC $\approx$ 90% Supervised Task-IL n/a n/a n/a
GSS [115] MLP $77.3\%\pm 0.5$ Supervised Task-IL n/a n/a n/a
HLOP-SNN [365] FC $95.15\%$ Supervised Task-IL,
Domain-IL
n/a n/a n/a
CIFAR-100
GPM [215] 5-layer
AlexNet
$72.48\%\pm 0.40$
(10-split)
Supervised Task-IL,
Class-IL
5.84M
parameters
770s (train) n/a
Adam-NSCL [218] ResNet-18 75.95% (20-split) Supervised Class-IL n/a n/a n/a
[223] ResNet18 $71.0\%\pm 0.3$ Supervised Task-IL n/a n/a n/a
GDumb [198] ResNet32 $60.3\%\pm 0.85$ OCL Task-IL $1105^{1}$ 60s (train) n/a
CoPE [192] Resnet18 21.62%±0.69 Supervised - $5000^{1}$
X-DER [199] ResNet18 49.93% Supervised Class-IL $\approx$ 50MB 5$\times$ more runtime
than DER++
n/a
ACAE-REMIND [210] Resnet-18 &
Resnet-32
62.30% AOC
(50 steps)
OCL Class-IL,
TA
12.8 MB n/a n/a
A-GEM [222] ResNet18 $\approx$ 62% Supervised Task-IL 10 times lower than EWC 100 times faster than EWC n/a
DVC [395] Resnet18 $24.1\%\pm 0.8$ OCL Class-IL $5000^{1}$ n/a n/a
SDAF [396] ResNet18 & MLP $39.0\%\pm 0.3$ OCL Class-IL $5000^{1}$ n/a n/a
SCR [193] ResNet18 & MLP $37.8\%\pm 0.3$ Supervised Class-IL $5000^{1}$approx 250s n/a
DSD-SNN [360] Multi-CONV &
1 FC layers
$77.92\%\pm 0.29$,
60.47% (10 steps)
Supervised Task-IL,
Class-IL
37.48% network compression n/a n/a
SOR-SNN [361] ResNet-18 $86.65\%\pm 0.20$
(20 steps)
Supervised Class-IL 0.32M
parameters
n/a n/a
HLOP-SNN [365] ResNet-18 78.58% Supervised Task-IL,
Domain-IL
n/a n/a n/a
MNIST
DER++ [197] FC $92.77\%\pm 1.05$ Supervised Domain-IL $500^{1}$ n/a n/a
EEC [205] 3-layer CONV
autoencoder
97.83% (10 classes) Supervised Class-IL 0.2MB disk space n/a n/a
CoPE [192] MLP $93.94\%\pm 0.20$ Supervised - $2000^{1}$ n/a n/a
GDumb [198] MLP $88.9\%\pm 0.6$ OCL Class-IL $300^{1}$ 60s (train) n/a
[294] 3 CONV layers $92.0\%\pm 0.1$ Unsupervised Task-IL n/a $2.5min^{2}$ n/a
DSD-SNN [360] Multi-CONV $97.30\%\pm 0.09$ Supervised Task-IL 34.38% network compression n/a n/a
Bayesian CL [304] Multi-FC layers $85.44\%\pm 0.16$ Supervised Class-IL 5× reduction n/a n/a
SpNCN [303] 4 FC layers 96.50% Supervised Task-IL n/a n/a n/a
ASP [102] 2 FC layers 94.20% Unsupervised Class-IL n/a n/a n/a
SpikeDyn[42] 2 FC layers Improved 23% (new task), 4% (old task) than ASP Unsupervised Class-IL ¡3000kB $0.2s^{3}$ 57% (train), 51%(inference) lower than ASP
TABLE XI: Continued.
lpspikecon